OpenStudy (anonymous):

If 400 feet of fencing is used to enclose a rectangular plot of land that borders a river, what is the maximum area that can be enclosed?

OpenStudy (anonymous):

It's pretty east to start out. Use the formulas: A = w * h P = 2h + w where P is the perimeter of the fence (400 ft), h is the length of one of the 2 parallel sides, and w is the length of the side that is parallel to the river.

OpenStudy (anonymous):

okay so 400 = 2h + w now what?

OpenStudy (anonymous):

system of equations?

OpenStudy (anonymous):

so solve for either h or w in that equation and substitute that into the Area formula

OpenStudy (anonymous):

okay so w = 200/h

OpenStudy (anonymous):

no quite

OpenStudy (anonymous):

so 400 = 2h + 200/h

OpenStudy (anonymous):

w = 400 - 2h

OpenStudy (anonymous):

argh okay ...

OpenStudy (anonymous):

so 400 = 2h + (400 - 2h)

OpenStudy (anonymous):

so that w should be substituted into the Area formula: A = w * h so A = (400 - 2h) * h

OpenStudy (anonymous):

A = 400 -2h^2

OpenStudy (anonymous):

A = 400h - 2h^2

OpenStudy (anonymous):

and now just find the derivative and set if equal to 0?

OpenStudy (anonymous):

and yes. take the derivative, set to 0, and that will give you h. you can find w from that, which will give you the total area

OpenStudy (anonymous):

so then h = 100 right?

OpenStudy (anonymous):

indeed it does!

OpenStudy (anonymous):

is w just the x coordinate at x =100?

OpenStudy (anonymous):

i mean y coordinate?

OpenStudy (anonymous):

sure. if you're mapping w (y-coord) against h (x-coord).

OpenStudy (anonymous):

okay so then my area would be 15000?

OpenStudy (anonymous):